# SOAC: Spatio-Temporal Overlap-Aware Multi-Sensor Calibration using Neural Radiance Fields

### 27 Nov 2023 v1

@@ Appendix E Qualitative results
In Figure 15 and Figure 16 are shown LiDAR/Camera projection on nuScenes and Pandaset sequences. The calibration optimized by SOAC provides substantially better alignment than the one from MOISST.
@@
Refer to caption Refer to caption Refer to caption
SOAC
@@
Refer to caption Refer to caption Refer to caption
SOAC
@@
Figure 15: More qualitative LiDAR/Camera reprojection results on nuScenes dataset.
@@
Refer to caption Refer to caption Refer to caption
SOAC
@@
Figure 16: Qualitative LiDAR/Camera reprojection results on Pandaset dataset.
Figure 17 shows the predicted images and masks from each NeRF trained with different cameras. The visibility masks are coherent with the predicted RGB images, allowing correct filtering for SOAC.
@@
Figure 17: Results using visibility grids on a Pandaset sequence – Prediction from different NeRFs.

### 27 Mar 2024 v3

@@ Appendix E Qualitative results
In Fig. 16 and Fig. 15 are shown LiDAR/Camera projection on nuScenes and Pandaset sequences. The calibration optimized by SOAC provides substantially better alignment than the one from MOISST.
@@
SOAC (ours)
Refer to caption Refer to caption Refer to caption
Figure 15: Qualitative LiDAR/Camera reprojection results on Pandaset [44] dataset.
@@
SOAC (ours)
Refer to caption Refer to caption Refer to caption
@@
SOAC (ours)
Refer to caption Refer to caption Refer to caption
Figure 16: More qualitative LiDAR/Camera reprojection results on nuScenes [3] dataset.
Fig. 17 shows the predicted images and masks from each NeRF trained with different cameras. The visibility masks are coherent with the predicted RGB images, allowing correct filtering for SOAC.
@@
Figure 17: Results using visibility grids on a Pandaset [44] sequence – Prediction from different NeRFs.